### v1

# Constant-Size Certificates for Leader Election in Chordal Graphs and Related Classes

Abstract
In distributed computing a certification scheme consists of a set of states and conditions over those states that enable each node of a graph to efficiently verify the correctness of a solution to a given problem. This work focuses on two fundamental problems: leader election and spanning tree construction. For each problem, we present a constant-size (per edge), local certification scheme, where the conditions available to each node can only refer to the graph induced by its one-hop neighborhood. In particular, we provide certification schemes for leader election in chordal and $K_{4}$-free dismantlable graphs and for spanning tree construction in dismantlable graphs, assuming a root is given. For chordal graphs, our leader election certification scheme additionally ensures an acyclic orientation, a property that is not generally verifiable using constant-size certificates in arbitrary graphs. To the best of our knowledge, these are the first local certification results tailored to these graph classes, potentially highlighting structural properties useful for verifying additional problems. Finally, we propose an algorithm that automatically transforms any certification scheme into a silent self-stabilizing algorithm (i.e., an algorithm that automatically recovers from faults) by adding only one extra state to the set of states of the certification scheme, assuming a Gouda fair scheduler. This transformation may be of independent interest.
keywords:
Local Certification, Distributed Verification, Proof Labelling Schemes, Chordal Graphs, Dismantlable Graphs, Leader Election, Distributed Computing
1 Introduction
Local certification [feuilloley2019introduction], also known as proof labelling schemes [korman2005proof], locally checkable proofs [goos2016locally] or distributed verification, verifies global properties through local conditions. Each node individually assesses these conditions and either accepts (i.e., all conditions are satisfied) or rejects (i.e., at least one condition is violated). An error is detected if any node rejects. For example, to verify that a graph is a cycle and not a path, each node accepts if its degree is two and rejects if its degree is one. However, the inverse problem of accepting a path and rejecting a cycle is not possible with certificates of constant size, even if the nodes are not identical [feuilloley2019introduction].
The ability to continue functioning and recovering from faults is essential for distributed systems. Dijkstra [dijkstra1974stabilising] introduced the concept of self-stabilization as a generic way to model various faults and guarantee the correct behavior of an algorithm regardless of transient faults. An algorithm is self-stabilizing if it eventually converges to a correct output from any arbitrary initial configuration (see also [dolev2000book]). A self-stabilizing algorithm is called silent if when it reaches a correct configuration, it remains in that configuration unless a fault occurs. On the other hand, a self-stabilizing algorithm that is not silent can transition between correct configurations as the algorithm perpetually continues computing. The design of a self-stabilizing algorithm typically consists of three components: (a) Detect when the global configuration is faulty (b) Reach a non-faulty configuration (c) Solve the original task. The first aspect naturally connects self-stabilization with local certification, where nodes maintain local certificates that allow verification of the global configuration via local checks. For instance, in the self-stabilizing spanning tree algorithm of [chen1991spanning], each node stores a pointer to its parent (as a way to solve the task) and its distance from the root (as a way to verify the solution). Without the latter, cycles cannot be locally detected.
As a form of fault-tolerance, verification is not the main task to be solved within a system and as a result it should not consume too many resources. Hence, in this work, we only consider certificates of constant size per edge. A significant challenge in this setting is that in general problems such as the detection of cycles mentioned previously require more than constant space. However, the existence of models such as Population Protocols [aspnes2009population] and Stone Age Computing [emek2013stoneage] which only use constant space demonstrates that such constraints do not preclude the existence of solutions to non-trivial problems.
The main problem that we consider in the following sections is Leader Election [LeLann77], which is an important primitive in distributed systems as it allows the coordination of processes in the system. The goal in this problem is for a unique node to enter and remain in the elected state indefinitely. A usual assumption within this context, that we do not make in this work, is that nodes have distinct identifiers. In the anonymous setting, the problem cannot be solved without additional assumptions due to symmetries [angluin1980local]. One way to overcome this impossibility is to characterize graph classes where the problem can be solved (e.g., as in [yamashita1996computing]) or to employ randomization (e.g., [itai1990symmetry]). In this work, we use the former method and provide certificates for leader election for the case of chordal and $K_{4}$-free dismantlable graphs.
Both chordal graphs and dismantlable graphs can be characterized by an ordering of their vertices. Chordal graphs admit an ordering called a simplicial elimination scheme, in which each vertex is adjacent to a clique of nodes that appear later in the ordering. Similarly, dismantlable graphs admit a dismantling order, where each vertex $v$ has a neighbor that is adjacent to all other neighbors of $v$ appearing later in the ordering. Dismantlability, in particular, is a property shared by many important classes in metric graph theory, such as bridged [FaJa, SoCh] and Helly graphs [BP-absolute, BaPr]. In this work, we design certificates that take advantage of these orderings to define and locally verify an orientation of the edges incident to each node, ensuring that the graph contains a unique sink node, which we define to be the leader.
1.1 Related Work
In the LOCAL model introduced by Linial [linial1987local, linial1992journal], it is assumed that nodes exchange messages with neighboring nodes until a task is solved. The complexity of an algorithm within this model thus depends on the size of the messages and on the number of synchronous communication rounds. An alternative definition of the model that is more appropriate for our setting, is to assume that every node can see the state of each node in its neighborhood up to some distance $d$ and based on this partial knowledge of the graph, choose an output. The goal in this case is to minimize $d$ and the size of states for some given task. What can be computed within this setting has received significant attention (e.g., [vishkin1986determinisitic, goldberg1987parallel, luby1986simple, alon1986fast]). The concept of locally verifiable problems was formalized by Naor and Stockmeyer in their seminal paper [naor1993lcl, naor1995journal] in which they defined the family of Locally Checkable Labeling (LCL) problems. A problem is in LCL if given the view, consisting of constant-sized input and output labels, at some constant distance $d$ for each node, there exists an algorithm that can decide whether a computed solution is correct. Although LCL problems were continually explored since their introduction, no groundbreaking results emerged until 2016 when Brandt et al. [brandt2016lower] led to a series of papers (e.g., [balliu2018classes, balliu2018locally, balliu2020randomness]) that gradually provided a complete picture on the classification of the round complexity of all LCL problems in four broad classes [suomela2020landscape].
Recently, Bousquet et al. [feuilloley2025certification] proposed a different direction of classification based on space instead of round complexity. Several results are given in [feuilloley2025certification], characterizing both “gaps” (i.e., pairs of functions $f(n),g(n)$, with $f(n)<g(n)$ where no property has optimal certificate size in $(f(n),g(n))$ for a given class of graphs) as well as proving the existence of properties with optimal certificate sizes for various classes of graphs including paths, cycles and caterpillars. In this work we consider chordal and ($K_{4}$-free) dismantlable graphs which are not explicitly studied in [feuilloley2025certification]. Our approach is complementary to that of [feuilloley2025certification] in the sense that we only consider specific problems but for larger classes of graphs.
The most similar paper to our work is [chalopin2024stabilising]. Therein, the authors propose a certificate and a set of conditions with constant memory per node, ensuring a unique sink for a simply connected subset of the regular triangular grid, where simply connected means that the configuration does not contain holes. Observe that simply connected parts of triangular grids are $K_{4}$-free dismantlable graphs, hence this work (in particular Section˜4) generalizes the result of [chalopin2024stabilising] and extends this idea to additional classes of graphs. Additionally, our definition of view is similar to the concept of binoculars introduced in [chalopin2015binoculars]. Binoculars provide each node $v$ with all information on the graph induced by the neighborhood of $v$ at distance one.
1.2 Our Contributions
We define constant-size certification schemes for verifying the existence of a unique sink in chordal graphs and $K_{4}$-free dismantlable graphs, under the assumption that each node can see the subgraph induced by its one-hop neighbors and the certificates of its neighbors. This notion of view can be seen as an intermediate assumption, requiring more information than standard radius one visibility but without accessing nodes at distance two. For chordal graphs, the certification scheme also ensures a valid acyclic orientation, which is particularly important as acyclicity cannot, in general, be verified with constant-size certificates. In the case of dismantlable graphs, we also provide a certification scheme to ensure the existence of a spanning tree, assuming there is a distinct node marked as the root. The certification scheme verifying the existence of a unique sink in $K_{4}$-free dismantlable graphs can be used to verify the existence of such a root. All of our certificates require only constant size per edge.
In the last section, we show that given a constant-size certificate, a silent self-stabilizing algorithm can always be constructed with constant memory per process, assuming a Gouda fair scheduler. Let $k$ be the number of bits needed to verify a task $T$ using a certificate. Our algorithm transforms the certificate to a self stabilizing algorithm using $k+1$ bits. This is less than prior general transformations, such as the one in [BlinFP14], which requires $O(k+\log n)$ bits, where $n$ is the number of computational entities in the system, but works even under an unfair scheduler. As a result, Algorithm 1 can be used in systems with constant memory where [BlinFP14] cannot be employed.
Our results demonstrate that self-stabilization with strong global guarantees (i.e., unique sink, acyclic orientation) is possible with limited local memory in natural graph classes. That is, we show that combining certificate-based verification and Gouda fairness provides a way for designing fault-tolerant systems even when only constant memory is available.
2 Model and Preliminaries
2.1 Model
Let $G=(V,E,\delta)$ be a graph with vertex set $V$, edge set $E$ and a port labeling function $\delta$. For any vertex $v\in V$, let $N(v)$ denote its neighborhood and let $N[v]=N(v)\cup\{v\}$ denote its closed neighborhood. The 1-ball at $v$, denoted by $B_{1}(v)$, is the graph induced by $N[v]$. When $G$ is oriented, we write $N^{+}(v)$ (resp. $N^{-}(v)$) for the outgoing (resp. incoming) neighborhood of $v$ in $G$ and refer to nodes in $N^{+}(v)$ as outneighbors of $v$. We define the port labeling function to be $\delta=\{\delta_{v}\}_{v\in V}$, where $\delta_{v}:N(v)\rightarrow\{1,2,\ldots,deg(v)\}$ assigns a unique port number to each edge incident to $v$ for every $v\in V$ (i.e., $\delta_{v}$ is a bijection). We write $K_{4}$ to denote a complete graph with four vertices. A graph is $K_{4}$-free if it contains no subgraph that is isomorphic to a $K_{4}$. We write directed triangle to refer to a cycle of size three.
Let $G=(V,E,\delta)$ be a graph and $S$ be an arbitrary set of states. For any $G$ and any mapping $x_{S}:V\rightarrow S$, we say that $(G,x_{S})$ is a configuration of $G$. We say that $x_{S}(v)$ is the state from $S$ assigned to $v\in V$ by $x_{S}$. We define a task $T=(O,P)$ as a set of states $O$ and a problem specification $P$, where $P$ is a set of configurations over $O$ that we call the correct configurations for $T$. It is possible to have several correct configurations in $P$ over the same graph. The notion of a correct configuration is specific to the task. In this paper, we consider two tasks: leader election (Sections˜3 and 4) and spanning tree construction (Section˜4). We give the definition of correct configurations in the problem specification $P$ and the set of states $O$ for each of these tasks here. We write $x_{O}$ to denote the mapping $x_{O}:V\rightarrow O$.
Leader Election
[…]
Spanning Tree
[…]
  • The set of directed edges $\{\vec{vw}\text{ }|\text{ }v\neq r\}$, where $w=\emph{parent}_{G}(v)$, is a directed spanning tree.
We slightly adapt the definition of local certification of [feuilloley2025certification] to our setting. We assume that the nodes of $G$ do not have identifiers, that is, $G$ is anonymous. Let $L$ be a possibly empty set of input labels. If $L\neq\emptyset$, we say that $G$ is labeled and we call the mapping from $V$ to $L$ the input function. Let $C$ be a non-empty set of states, called the certificate. A certificate assignment on $G$ is a mapping $x_{C}:V\rightarrow C$. For a given certificate assignment $x_{C}$ and a node $v\in V$, the view of $v$ consists of the 1-ball $B_{1}(v)$ and the restriction of the port numbering function $\delta$, the input function (if defined) and the certificate assignment $x_{C}$ to $B_{1}(v)$. An example of the view of a node is given in Figure˜1.
Figure 1: Right: A graph with port labels (red) where each node is assigned a certificate in $\{0,1\}^{*}$ (black). Left: The view of the node marked in gray.
Let $R$ be a set of configurations of the form $(B,x_{C})$ where $B$ is a ball of radius 1 and $x_{C}$ is a certificate assignment defined on the nodes of $B$. We refer to $R$ as a set of local configurations. A local configuration is said to be correct with respect to $(C,R)$ if for every $v\in V$ the pair $(B_{1}(v),x_{C}|_{N[v]})$ belongs to $R$, where $x_{C}|_{N[v]}$ denotes the restriction of $x_{C}$ to $N[v]$. In other words, the local configuration is correct if the labeling in each ball of radius 1 satisfies the local conditions defining a correct configuration $R$. We refer to the pair $(C,R)$ as a certification scheme.
A certification scheme is used to solve a given task $\mathcal{T}=(O,P)$. The key property we prove, is that if a configuration is locally correct with respect to the certification scheme, then it is globally correct (i.e., it is a correct configuration for the task). Formally, let $out:C\rightarrow O$. We say that there exists a certification scheme $(C,R)$ for some task $(O,P)$ in a family of possibly labelled graphs $\mathcal{F}$ if for every $G\in\mathcal{F}$ the following two properties are satisfied [^1] The symbol $\circ$ denotes the composition operator.:
Existence
$\forall G\in\mathcal{F},\text{ }\exists(G,x_{C})\text{ s.t. }\forall v\in V,\text{ }(B_{1}(v),x_{C}|_{N[v]})\in R$
Soundness
$\forall(G,x_{c})\text{ s.t. }G\in\mathcal{F},\text{ if }\forall v\text{ }(B_{1}(v),x_{C}|_{N[v]})\in R\text{ then }(G,\emph{out}\circ x_{C})\in P$
In most of our certification schemes, the certificates correspond to an edge orientation. For these certificates, we denote the configuration as $\vec{G}$ and we adopt the following encoding for the certificates: the certificate assigned to each node $v\in V$ is composed of $\deg(v)$ bits where each bit encodes the direction of the corresponding incident edge. We assume that 0 denotes an incoming edge to $v$ and 1 denotes an outgoing edge from $v$. The bit at position $i\in\{0,\ldots,\deg(v)-1\}$ of the certificate corresponds to the direction of the edge at port $i$ of $v$. A visualization of this for the graph in Figure˜1 is given in Figure˜2.
Figure 2: The $i$-th bit of the certificate (black numbers) of each node $v$ corresponds to the direction of the edge incident to port $i$ (red numbers) of $v$, for $i\in\{0,\ldots,deg(v)-1\}$ where 0 denotes an incoming and 1 an outgoing edge. On the left we show the direction of the edges and on the right we additionally show the bit of the certificate of each node corresponding to each edge.
For each edge $uv$ where $u$ (resp. $v$) is connected to $v$ (resp. $u$) through port $i$ (resp. $j$), we say that the edge is oriented if the $i$-th bit of $u$ differs from the $j$-th bit of $v$.
2.2 Warm Up: A Certificate for Leader Election in Trees
We give a simple certificate that guarantees a unique sink in the class of trees. While not new (e.g., one could easily construct a certificate following the self–stabilizing leader election algorithm of [antonoiu1996trees]), we believe it gives a good intuition behind the certificates and proofs in the following sections. The certificate of each node consists of the orientation of edges incident to the node. A correct configuration is defined by two conditions, T1 and T2. If both T1,T2 hold at every node of a tree $T$, then $T$ has a unique sink which we define to be the leader.
  1. T1
    Each edge is directed.
  2. T2
    For each vertex $v$, $v$ has at most one outgoing edge.
An orientation satisfying $T_{1}$ and $T_{2}$ (i.e., satisfying the Existence property) exists: first direct each edge incident to a leaf toward its unique neighbor, then iteratively orient edges away from leaves in the subgraph of remaining unoriented edges.
Theorem 2.1 (Soundness).
In an unlabeled tree $T$, any configuration $\vec{T}$ satisfying T1 and T2 contains a unique sink.
Proof 2.2.
We prove the theorem by induction on the number of vertices of $T$. The base case of a single vertex is trivial. Suppose $T$ contains at least two vertices and let $v$ be a leaf of $T$. Let $\vec{T}$ be a configuration of $T$ and assume that $\vec{T}$ satisfies T1 and T2. Observe that the configuration $\vec{T}\setminus\{v\}$ of $T\setminus\{v\}$ also satisfies T1 and T2. By the induction hypothesis $\vec{T}\setminus\{v\}$ has a unique sink $t$. We show that $\vec{T}$ also has a unique sink.
Call $u$ the neighbor of $v$ in $T$. Assume $vu$ is oriented $\overrightarrow{uv}$. Then from T2, $u$ has no other outneighbor in $\vec{T}$, so $u$ is the unique sink in $\vec{T}\setminus\{v\}$. Since $u$ has an outgoing edge in $\vec{T}$, it is not a sink, but $v$ is. Thus, $v$ is the unique sink in $\vec{T}$. Assume now $vu$ is oriented $\overrightarrow{vu}$. Then $t$ remains a sink in $\vec{T}$ and $v$ is not a sink, so $t$ is the unique sink. Hence, in all cases $\vec{T}$ has a unique sink, completing the proof.
3 Chordal Graphs
We generalize the certification scheme from Section 2.2 to ensure the validity of a certificate assignment, which guarantees acyclicity and a unique sink in chordal graphs. Each node’s certificate is the orientation of its incident edges. We refer to the nodes of $G$ and the certificate assignment as a configuration $\vec{G}$. A graph is chordal if every induced cycle is of length three. Every chordal graph has a simplicial vertex $v$, that is, a vertex $v$ such that $N(v)$ is a clique. We define a configuration to be correct for leader election, if for every vertex $v$ of a chordal graph, the following local conditions hold:
  1. C1
    Each edge incident to $v$ is directed.
  2. C2
    All outneighbors of $v$ are adjacent, that is, $N^{+}(v)$ induces a complete graph.
  3. C3
    Vertex $v$ does not belong to any directed triangle.
Observe that the verification of C3 is possible, since the view by definition contains the orientation of edges between neighboring nodes.
{observation}
[Existence] A correct configuration with respect to C1–C3 exists and can be constructed as follows: direct all edges incident to a simplicial vertex $u$ toward $N(u)$, then iteratively restrict the graph to the subgraph of remaining unoriented edges and orient edges away from simplicial vertices.
Lemma 3.1.
In a chordal graph $G$, any configuration $\vec{G}$ satisfying C1–C3 is an acyclic orientation of $G$.
Proof 3.2.
Suppose the lemma statement does not hold and $\vec{G}$ contains an oriented cycle. Take $\mathsf{C}=(u_{0},u_{1},\ldots,u_{k-1})$ to be the smallest such cycle. From C3, $k>3$. Then without loss of generality, $u_{1}$ is adjacent to at least three edges: $\overrightarrow{u_{0}u_{1}}$, $\overrightarrow{u_{1}u_{2}}$ and some directed edge $u_{1}u_{i}$ for $3\leq i<k$. Depending on the orientation of $u_{1}u_{i}$, $G$ contains either the oriented cycle $\mathsf{C}^{\prime}=(u_{1},u_{i},u_{i+1},\ldots,u_{k-1},u_{0})$ or the oriented cycle $\mathsf{C}^{\prime\prime}=(u_{1},u_{2},\ldots,u_{i},u_{1})$. In either case, a smaller cycle than $\mathsf{C}$ exists in $G$, contradiction.
Lemma 3.3 (Soundness).
In a chordal graph $G$, any configuration $\vec{G}$ satisfying C1–C3 contains a unique sink.
Proof 3.4.
We prove the lemma by induction on the number of vertices of $G$. The base case holds for $G$ having one vertex. For $G$ with at least two vertices, let $u$ be a simplicial vertex, that is, $N(u)$ induces a complete subgraph. Consider a configuration $\vec{G}$ of $G$ satisfying C1–C3. Removing $u$ from $\vec{G}$ gives a configuration of $G\setminus\{u\}$ that also satisfies C1–C3, so by the induction hypothesis $\vec{G}\setminus\{u\}$ has a unique sink $t$.
We claim that if $u$ has an outneighbor $v$ in $\vec{G}$, then $t$ is also the unique sink in $\vec{G}$. Due to $\vec{G}\setminus\{u\}$ having a unique sink, no vertex other than $t$ can be a sink in $\vec{G}$. If $t$ is not a sink in $\vec{G}$, then $\overrightarrow{tu}\in\vec{G}$, as all remaining edges incident to $t$ are incoming due to $t$ being a sink in $\vec{G}\setminus\{u\}$. Since $v,t\in N(u)$ and $u$ is a simplicial vertex, $v\in N^{-}(t)$. Thus $\overrightarrow{tu},\overrightarrow{uv},\overrightarrow{vt}$ induce a directed triangle, contradicting C3.
Suppose now that $u$ has no outneighbor, (i.e., $u$ is a sink of $\vec{G}$). For any neighbor $v$ of $u$, any outneighbor $w$ of $v$ in $\vec{G}\setminus\{u\}$ must be a neighbor of $u$ in $\vec{G}$ from C2. Consequently, $N^{+}(v)\subseteq N[u]$. We claim that there exists a vertex $t^{\prime}\in N(u)$ that has no outneighbor in $N(u)$. Suppose otherwise. Then, every $v\in N(u)$ has an outneighbor in $N(u)$ and $\vec{G}\setminus\{u\}$ contains a cycle, which is a contradiction by Lemma 3.1. Since $N(u)$ is a complete graph, this $t^{\prime}$ is unique. Furthermore, as we have already shown that $t^{\prime}$ has no outneighbor outside of $N[u]$, $t^{\prime}$ is the unique sink of $\vec{G}\setminus\{u\}$. Since $^{\prime}\in N^{-}(u), tt$ is not a sink in $\vec{G}$. Therefore, $u$ is the unique sink of $\vec{G}$.
Theorem 3.5 directly follows from Section˜3 and Lemmas˜3.1 and 3.3.
Theorem 3.5.
In a chordal graph $G$, any configuration $\vec{G}$ satisfying C1–C3 is an acyclic orientation of $G$ containing a unique sink.
4 Dismantlable graphs
This section is divided into two parts: a certification scheme ensuring a unique sink in $K_{4}$-free dismantlable graphs and a certification scheme for a rooted spanning tree, given a unique root, in dismantlable graphs. Let $G$ be a graph and $u,v$ be two vertices in $V(G)$. We say that $v$ is dominated by $u$ in $G$ if $N[v]\subseteq N[u]$. A graph is dismantlable if there exists an ordering $v_{1},\ldots,v_{n}$ of $V(G)$ such that in $G[\{v_{1},\ldots,v_{i}\}]$, $v_{i}$ is dominated by some $v_{j}$ with $j<i$. So for each $i>1$, there exists $j<i$ such that $N[v_{i}]\cap\{v_{1},\ldots,v_{i}\}\subseteq N[v_{j}]$. Dismantlable graphs are exactly the class of cop-win graphs in the classic Cops and Robbers game [aigner1984cops, nowakowski1983vertex, quilliot1978jeux] and include well-structured families such as Bridged graphs and Helly graphs, which are known for their convexity and metric properties.
4.1 Leader election in $K_{4}$-free dismantlable graphs
In this section we only consider $K_{4}$-free dismantlable graphs. As in Section 3, the certificate at each node consists of an orientation on the edges incident to the node and we write $\vec{G}$ to denote a configuration. We define a locally correct configuration to be one that satisfies the following rules for every node $v$ of $G$. We show that if all rules are satisfied, then $G$ contains a unique sink that we define to be the leader.
  1. D1
    Each edge incident to $v$ is directed.
  2. D2
    Either $v$ has no outgoing edges, or there exists an outneighbor $v^{+}$ of $v$ adjacent to all other outneighbors of $v$, i.e, $N^{+}(v)\subseteq N[v^{+}]$.
  3. D3
    Vertex $v$ does not belong to any directed triangle.
{observation}
[Existence] Any dismantlable graph admits an acyclic orientation satisfying D1–D3. {claimproof} Let $v_{1},v_{2}\ldots,v_{n}$ be the dismantling order of the nodes of some $K_{4}$-free dismantlable graph $G$. Orient each edge between two nodes $v_{i}$ and $v_{j}$ as $\overrightarrow{v_{i}v_{j}}$ if and only if $i>j$ in the dismantling order. This orientation satisfies D1 and D3. For D2, consider any node $v$. If $N^{+}(v)=\emptyset$, D2 is immediately satisfied for $v$. So suppose $N^{+}(v)\neq\emptyset$. By dismantlability, every node $v$ is dominated by some later node $u$ in the dismantling order. Then $\overrightarrow{vu}$ is an edge of the orientation and for every other outneighbor $w$ of $v$, $w\in N(u)$. Hence, $N^{+}(v)\subseteq N[u]$, satisfying D2.
However, notice that in the case of $K_{4}$-free dismantlable graphs, one cannot hope to have a property similar to Lemma˜3.1. Indeed, as in [chalopin2024stabilising], one can construct configurations satisfying D1-D3 that contain directed cycles. An example is shown in Figure˜3.
Figure 3: A $K_{4}$-free dismantlable graph and a certificate assignment that satisfies D1–D3 but is not acyclic. The square node dominates all cyclic nodes. The edges forming the cycle are marked in orange.
Theorem 4.1 (Soundness).
Let $G$ be a $K_{4}$-free dismantlable graph. If a configuration $\vec{G}$ satisfies D1–D3, then $\vec{G}$ contains a unique sink.
Proof 4.2.
Consider a graph $G$ and a configuration $\vec{G}$ of $G$ satisfying conditions D1–D3. We prove the theorem by induction on the number of vertices in $G$. The theorem trivially holds if $G$ has one vertex. Suppose $G$ has at least two vertices. Let $v$ be a vertex dominated by some vertex $u$ in $G$, and let $G^{\prime}=G\setminus\{v\}$ and $\vec{G}^{\prime}=\vec{G}\setminus\{v\}$. Since D1 and D3 hold in $\vec{G}$, they also hold in $\vec{G}^{\prime}$. To apply the induction hypothesis to $\vec{G}^{\prime}$, it remains to show that D2 holds in $\vec{G}^{\prime}$. For any vertex $w$ such that $v\notin N^{+}(w)$ in $\vec{G}$, condition D2 remains satisfied for $w$ in $\vec{G}^{\prime}$. More precisely, if D2 is not satisfied for $w$ in $\vec{G}^{\prime}$, then $v\in N^{+}(w)$ and $v$ is the only possible choice for $w^{+}$ in $\vec{G}$. We first establish a useful lemma.
Lemma 4.3.
For any vertex $w\in N^{-}(v)\cap N^{-}(u)$, we can assume that $w^{+}\neq v$.
Proof 4.4.
If $w^{+}=v$, i.e., $N^{+}(w)\subseteq N[v]$, then $N^{+}(w)\subseteq N[v]\subseteq N[u]$ and we can thus choose $w^{+}=u$.
In the remainder of the proof we show that D2 is satisfied in $\vec{G}^{\prime}$ and that $\vec{G}$ contains a unique sink. We distinguish two cases, depending on whether $uv$ is directed as $\overrightarrow{uv}$ or $\overrightarrow{vu}$.
Case 1.
The edge $uv$ is $\overrightarrow{vu}$.
{claimproof}
Consider a vertex $w\neq v$ such that $w\in N(v)$. Observe that if $w\in N^{+}(v)$ and D2 is satisfied in $\vec{G}$ for $w$ then D2 cannot be broken for $w$ in $\vec{G}^{\prime}$. So suppose that $w\in N^{-}(v)$. Since $u\in N^{+}(v)$, by D3, $w\in N^{-}(u)$. Consequently, by Lemma˜4.3, we can assume that $w^{+}\neq v$ and D2 is thus satisfied for $w$ in $\vec{G}^{\prime}$.
Since D1–D3 are satisfied in $\vec{G}^{\prime}$, $\vec{G}^{\prime}$ contains a unique sink $t$. Observe that if $t\in N^{-}(v)$ in $\vec{G}$, $\overrightarrow{tv},\overrightarrow{vu},\overrightarrow{ut}$ induce a directed cycle in $\vec{G}$, a contradiction. Consequently, $t$ is also a sink in $\vec{G}$. Since all other vertices of $\vec{G}^{\prime}$ have outgoing edges in $\vec{G}^{\prime}$ (and thus in $\vec{G}$), and since $v$ is not a sink in $\vec{G}$ (as $v\in N^{-}(u)$), $t$ is the unique sink of $\vec{G}$. This ends the proof of ˜1.
Case 2.
The edge $uv$ is $\overrightarrow{uv}$.
{claimproof}
We further distinguish two cases, depending on whether we can assume that $u^{+}=v$.
Case 2.1.
$N^{+}(u)\subseteq N[v]$, that is, we can assume $u^{+}=v$.
{claimproof}
Since $v$ is dominated by $u$, for any $w\in N(v)$ we also have $w\in N(u)$. If additionally $w\in N^{-}(u)$, then by D3 it follows that $w\in N^{-}(v)$. Consequently, for any $w\in N(v)$, it must be that $w\in N^{+}(u)\cap N^{+}(v)$, or $w\in N^{-}(u)\cap N^{-}(v)$, or $w\in N^{+}(u)\cap N^{-}(v)$.
We distinguish two cases, depending on whether $N^{+}(u)\cap N^{-}(v)$ is empty.
Case 2.1.1.
$N^{+}(u)\cap N^{-}(v)=\emptyset$
{claimproof}
We first show that D2 is satisfied in $\vec{G}^{\prime}$. Consider a vertex $w\neq v$ such that $w\in N^{-}(v)$. If $w\neq u$, since $u$ dominates $v$, $u$ must be a neighbor of $w$. Given that $N^{+}(u)\cap N^{-}(v)=\emptyset$, it follows that $w\in N^{-}(u)$. By Lemma˜4.3, we can assume $w^{+}\neq v$, thus D2 is satisfied for $w$ in $\vec{G}^{\prime}$.
We now show that D2 holds for $u$. Since $N^{+}(u)\subseteq N[v]$ and $N^{+}(u)\cap N^{-}(v)=\emptyset$, we have $N^{+}(u)\subseteq N^{+}(v)\cup\{v\}$. Given that $N^{+}(v)\subseteq N(v)\subseteq N[u]$ and $v\in N^{+}(u)$, by D3, we obtain $N^{+}(v)\subseteq N^{+}(u)$. Therefore, $N^{+}(u)=N^{+}(v)\cup\{v\}$. Hence, either $N^{+}(v)=\emptyset$ and $N^{+}(u)=\{v\}$, or there exists $v^{+}\in N^{+}(v)$ such that $N^{+}(v)\subseteq N[v^{+}]$, implying $N^{+}(u)=N^{+}(v)\cup\{v\}\subseteq N[v^{+}]$. In either case, D2 is satisfied for $u$ in $\vec{G}^{\prime}$.
By induction hypothesis $\vec{G}^{\prime}$ contains a unique sink $t$. Notice that as $N^{+}(u)=N^{+}[v]$, $v$ is a sink of $\vec{G}$ if and only if $u$ is a sink of $\vec{G}^{\prime}$. Therefore, if $t\neq u$ in $\vec{G}^{\prime}$, $v$ is not a sink in $\vec{G}$ and $t$ is the unique sink in $\vec{G}$. If $t=u$, then $t$ is not a sink in $\vec{G}$ but $v$ is, consequently $v$ is the unique sink of $\vec{G}^{\prime}$. In either case, $\vec{G}$ contains a unique sink concluding the proof of ˜2.1.1.
Case 2.1.2.
$N^{+}(u)\cap N^{-}(v)\neq\emptyset$
{claimproof}
In this case, we consider the configuration $\vec{G}^{*}$ obtained from $\vec{G}$ where for each $w\in N^{+}(u)\cap N^{-}(v)$ in $\vec{G}$, we have replaced the arc $\overrightarrow{uw}$ by the arc $\overrightarrow{wu}$. We show that D1–D3 are satisfied in $\vec{G}^{*}$. Observe that since all edges are directed in $\vec{G}$, they are all directed in $\vec{G}^{*}$ and D1 holds.
Suppose that there exists a directed triangle in $\vec{G}^{*}$. Then necessarily, this triangle contains an edge $\overrightarrow{wu}$ with $w\in N^{+}(u)\cap N^{-}(v)$ in $\vec{G}$. Let $x$ be the vertex such that $\overrightarrow{wu}$, $\overrightarrow{ux}$, $\overrightarrow{xw}$ form a triangle in $\vec{G}^{*}$. Note that $x\neq v$ as $w\in N^{-}(v)\cap N^{+}(x)$ in $\vec{G}$ and $\vec{G}^{*}$. Since $\overrightarrow{ux}$ is in $\vec{G}^{*}$, $x\in N^{+}(u)$ in $\vec{G}^{*}$ and in $\vec{G}$ (the set of outneighbors of $u$ in $\vec{G}^{*}$ is a subset of its set of outneighbors in $\vec{G}$). Since in $\vec{G}$, $N^{+}(u)\subseteq N[v]$, $x$ is a neighbor of $v$. Since $w\in N^{+}(x)$ and $v\in N^{+}(w)$, by D3, we have $x\in N^{-}(v)$ in $\vec{G}$. Consequently, in $\vec{G}$, $x\in N^{+}(u)\cap N^{-}(v)$, and thus in $\vec{G}^{*}$, $x\in N^{-}(u)$, a contradiction. This shows that D3 holds in $\vec{G}^{*}$.
We now show that D2 holds in $\vec{G}^{*}$. For any $\notin(N^{+}(u)\cap N^{-}(v))\cup\{u\}, xx$ has the same set of outneighbors in $\vec{G}$ and in $\vec{G}^{*}$, and thus D2 holds for $x$ in $\vec{G}^{*}$ as it holds in $\vec{G}$. Since $N^{+}(u)\subseteq N(v)$ in $\vec{G}$, since $N^{+}(u)$ in $\vec{G}^{*}$ is a subset of $N^{+}(u)$ in $\vec{G}$, and since $v\in N^{+}(u)$ in $\vec{G}^{*}$, D2 also holds for $u$ in $\vec{G}^{*}$ (with $u^{+}=v$). Consider now $w\in(N^{+}(u)\cap N^{-}(v))$ and let $w^{+}\in N^{+}(w)$ such that $N^{+}(w)\subseteq N[w^{+}]$ in $\vec{G}^{\prime}$. As $u\notin N^{+}(w)$ in $\vec{G}$, $w^{+}\neq u$. Since $v\in N^{+}(w)$, $w^{+}\in N[v]\setminus\{u\}\subseteq N(u)$. Hence, $w^{+}\in N(u)$, implying $u\in N[w^{+}]$ and thus in $\vec{G}^{*}$ we also have $N^{+}(w)\subseteq N[w^{+}]$. Therefore, D2 is satisfied at $w$ in $\vec{G}^{*}$. Hence D2 holds in $\vec{G}^{*}$.
Consequently, by ˜2.1.1 applied to $\vec{G}^{*}$, we deduce that there exists a unique sink $t$ in $\vec{G}^{*}$. Note that $t\notin(N^{+}(u)\cap N^{-}(v))\cup\{u\}$ as every vertex $w\in(N^{+}(u)\cap N^{-}(v))\cup\{u\}$ has $v$ as an outneighbor in $\vec{G}$ and in $\vec{G}^{*}$. So, $t$ is a sink in $\vec{G}$ and since no node of $(N^{+}(u)\cap N^{-}(v))\cup\{u\}$ is a sink in $\vec{G}$, $t$ is the unique sink of $\vec{G}$. This ends the proof of ˜2.1.2. This ends the proof of ˜2.1.
Case 2.2.
$N^{+}(u)\not\subset N[v]$, implying that $u^{+}\neq v$.
{claimproof}
We first show that D2 is satisfied in $\vec{G}^{\prime}$. For D2 to be violated, $v$ should be the only choice for $w^{+}$, for some $w\neq u$. If $N^{+}(w)=\{v\}$, D2 trivially holds for $w$ in $\vec{G}^{\prime}$. If $u\in N^{+}(w)$, D2 holds for $w$ by Lemma˜4.3. Suppose now there exists $z\in N^{+}(w)\setminus\{v\}$. Then $z\in N(v)$ and from $u$ dominating $v$, $\{w,z\}\in N(u)$. Hence, $v,u,w,z$ induce a $K_{4}$, contradiction. So we can assume that $w^{+}\neq v$ for all $w\in V$, and D2 holds for all nodes in $\vec{G}^{\prime}$.
By induction hypothesis, $\vec{G}^{\prime}$ contains a unique sink $t$. We consider two cases, based on the orientation of $u^{+}v$.
Case 2.2.1.
The edge $u^{+}v$ is $\overrightarrow{vu^{+}}$, that is, $u^{+}\in N^{+}(v)$.
{claimproof}
If $t$ is not a sink in $\vec{G}$, $t\in N^{-}(v)$ and thus, $t\in N(u)$. Observe that it cannot be that $t=u^{+}$ since we have assumed $\overrightarrow{vu^{+}}$, which means that $vt$ is oriented as $\overrightarrow{vt}$ and $t$ would still be a sink in $\vec{G}$ contradicting our assumption. So suppose $t\neq u^{+}$. Since $t$ is a sink in $\vec{G}^{\prime}$, $t\in N^{+}(u)$ and thus $t\in N(u^{+})$, forming a $K_{4}$ among $u,v,t,u^{+}$. Therefore, $t\notin N^{-}(v)$ and $t$ is also a sink in $\vec{G}$. Since $v$ is not a sink in $\vec{G}$ due to $\overrightarrow{vu^{+}}$, $t$ is the unique sink in $\vec{G}$. This ends the proof of ˜2.2.1.
Case 2.2.2.
The edge $u^{+}v$ is $\overrightarrow{u^{+}v}$, that is, $u^{+}\in N^{-}(v)$.
{claimproof}
We claim that $u^{+}=t$. Suppose for contradiction that $u^{+}\neq t$. Then, since D2 is satisfied in $\vec{G}$, there exists $u^{++}\in N^{+}(u^{+})$ such that $N^{+}(u^{+})\subseteq N[u^{++}]$. Since we have assumed that $v\in N^{+}(u^{+})$, it must be that $v\in N[u^{++}]$. Suppose $v\neq u^{++}$, that is, $v\in N(u^{++})$. Then since $v$ is dominated by $u$, the node $u^{++}\in N(u)$. However, this means that the vertices $v,u,u^{+}$ and $u^{++}$ induce a $K_{4}$, hence this case is not possible. So suppose $v=u^{++}$. Suppose that $u^{+}$ has some outneighbor $w$ distinct from $v$. Then, since $v=u^{++}$, $w$ must be in $N(v)$ and $u$ is connected to $w$ as $u$ dominates $v$. In this case, the vertices $v,u,u^{+}$ and $w$ induce a $K_{4}$ so this setting is not possible and it must be that $N^{+}(u^{+})=\{u^{++}\}=\{v\}$. Therefore, it must be that $u^{+}=t$ in $\vec{G}^{\prime}$, contradicting the initial assumption.
Consider now $u^{+}=t$ in $\vec{G}^{\prime}$, implying that $v$ is the only outneighbor of $u^{+}$ in $\vec{G}$. Let $z\in N(v)$. Then, $z$ must also be neighboring to $u$ as $u$ dominates $v$. We claim that $v$ is the unique sink in $\vec{G}$. The orientation $\overrightarrow{uz}$ means that the edge $u^{+}z$ must also exist implying that $u,v,u^{+}$ and $z$ induce a $K_{4}$ which is not possible. So $uz$ must be oriented as $\overrightarrow{zu}$. From D3, $zv$ should be oriented as $\overrightarrow{zv}$. Hence every neighbor of $v$ in $\vec{G}$ must be incoming to $v$. Therefore, $v$ is a sink in $\vec{G}$ and since the unique sink $u^{+}$ of $\vec{G}^{\prime}$ is not a sink in $\vec{G}$ due to $\overrightarrow{u^{+}v}$, $v$ is the unique sink in $\vec{G}$. This ends the proof of ˜2.2.2. This ends the proof of ˜2.2. This ends the proof of ˜2. Theorem 4.1 follows from the analysis in Cases 1 and 2.
4.2 Spanning Tree
We present a certificate verifying the existence of a spanning tree in any dismantlable labeled graph $G=(V,E)$. The labeling function assigns each $v\in V$ a label in $L=\{0,1\}$, with exactly one node labeled 1 (the root $r$) and all others labeled 0. We define $C=\{0,1,2\}$ to be the certificate and each node is assigned a value from $C$ by $x_{C}:V\rightarrow C$. The following must hold for every node $v$ of $G$ with value $i$.
  1. N1
    Node $v\neq r$ has at least one neighbor with value $(i-1)\mod 3$.
  2. N2
    Node $v$ is not in any three-node triangle labeled $\{0,1,2\}$.
We show that if N1–N2 are satisfied for every node in $G$, adding an edge from any node $v$ distinct from the root with value $i$ to any neighbor of $v$ with value $i-1\mod 3$, induces a spanning tree.
Lemma 4.5 (Existence).
In any graph $G$ there exists a certificate assignment that satisfies N1–N2 for every node in $G$.
Proof 4.6.
We write $d(v,r)$ to denote the distance between a node $v$ and the root $r$. We construct the assignment as follows. For each node $v$ at distance $k>1$ from the root in $G$, assign $d(v,r)\mod 3$ to $v$. This satisfies N1. Let $u,w$ be two neighbors of $v$ that are also neighboring to each other. Without loss of generality assume that $v$ is the closest vertex to $r$ among the triangle $u,v,w$ and let $k=d(v,r)$. Then $k\leq d(u,r)\leq k+1$ and $k\leq d(w,r)\leq k+1$, thus N2 is satisfied and an assignment that satisfies N1–N2 exists.
[…]
Lemma 4.7.
If N1–N2 are satisfied for every node in a dismantlable graph $G$, $\vec{G}$ does not contain oriented cycles.
Proof 4.8.
Suppose the lemma does not hold. We define the order of a vertex $v$ as its position in the dismantling order. Let the sum of orders of a cycle be the sum of the orders of its vertices. Among all directed cycles in $\vec{G}$, let $\mathsf{C}=(v_{1},v_{2},\ldots,v_{k})$ be the cycle that minimizes this sum. Notice that the values of all nodes in $\mathsf{C}$ must follow the sequence $\ldots,0,1,2,0,1,2,\ldots$ so that the orientation of the cycle remains consistent. Call $v$ the node with the largest order in $\mathsf{C}$, $u$ the node dominating $v$ and let $w,z$ be the two neighbors of $v$ in $\mathsf{C}$. Observe that $u$ is also adjacent to $w,z$ and that by definition its order is smaller than the one of $v$. Without loss of generality let the value of $v$ be 1 and let the values of $w,z$ be 0 and 2 respectively. Then, the value of $u$ must be 1, otherwise one of $(u,v,w)$ and $(u,v,z)$ breaks N2. We show that there exists a cycle $\mathsf{C}^{\prime}$ with a smaller sum of orders than $\mathsf{C}$ which includes $u$ instead of $v$. First suppose that $u\in\mathsf{C}$ and let $u_{-}$ be the predecessor of $u$ in $\mathsf{C}$. Then $\mathsf{C}^{\prime}$ is induced by the nodes in $\mathsf{C}\setminus\{v,z,\ldots,u_{-}\}$. Suppose now that $u\notin\mathsf{C}$ and let $\mathsf{C}^{\prime}$ be the cycle induced by the nodes in $\{u\}\cup\mathsf{C}\setminus\{v\}$. In either case, the existence of $\mathsf{C}^{\prime}$ contradicts $\mathsf{C}$ being a cycle with minimum sum of orders, proving the lemma.
Let $T$ be a spanning subgraph of $\vec{G}$ in which each node $v$, except the root, arbitrarily selects a parent among its neighbors with value equal to $(\textit{value of }v-1)\mod 3$. We show that $T$ is a tree. Since $\vec{G}$ is acyclic by Lemma 4.7, it follows that $T$ is also acyclic. We prove that $T$ is connected.
Lemma 4.9.
If N1–N2 are satisfied for every node in a dismantlable graph, $T$ is connected.
Proof 4.10.
Let $u$ be a node in $T$. Since from Lemma 4.7, $\vec{G}$ and as a result $T$ does not contain cycles, the directed path starting at $u$ must end at a node without a successor in $T$. From N1, the root is the only node without a smaller neighbor. Hence, any node $u$ must be in a path that reaches the root, and as a result $T$ is connected.
Finally, since every node in $G$ has a label and must be in $T$, from Lemmas 4.7 and 4.9 we obtain the following:
Theorem 4.11 (Soundness).
In a dismantlable graph, if N1–N2 are satisfied for every node, $T$ is a spanning tree.
5 From Certificate to Self–Stabilization
An algorithm is called self-stabilizing when it can autonomously recover from transient faults. In particular, a distributed self-stabilizing algorithm is usually defined as a set of rules evaluated at every node. Each rule is of the form:
$\texttt{name-of-rule}:\langle guard\rangle\longrightarrow command$
where guard is a boolean predicate over the variables of the closed neighborhood of the node and the command specifies how the node updates its variables. Here, we assume that each node only updates its state variable. A self-stabilizing algorithm is said to be silent if after some finite time no node updates its variables anymore. When activated, each node can perform computations based on its state and the states of its neighbors and update its own state. The order of node activations is determined by a scheduler. A scheduler is characterized both by synchronicity and fairness. In particular, here we consider an asynchronous scheduler meaning that at each step, an arbitrary (non-empty) subset of nodes may be activated. In terms of fairness, we assume a Gouda fair scheduler [Gouda2001theory]. According to [dubois2011taxonomy], Gouda fairness is the strongest fairness assumption. It requires that for any configuration that appears infinitely often during the execution of an algorithm, every configuration that is reachable by executing any non-empty subset of enabled commands (i.e., commands for which the guard is true) also appears infinitely often. We assume that the graphs we consider are finite. In this setting, Gouda fairness ensures that any configuration that is infinitely often reachable is eventually reached. In this section we show how to obtain a silent self–stabilizing algorithm that works under a Gouda fair scheduler, given a certification scheme.
We say that under a given algorithm $A$, a certificate assignment is final if no node satisfies any of the guards in $A$. We define a correct final certificate assignment to be one that is both final with respect to an algorithm $A$ and correct with respect to a certification scheme. Call the set of states in a given certificate $C$. Informally, every node that detects that a condition of the certification scheme is violated sets its state to a reset state that we denote $\bot$. Additionally, every node that has at least one neighbor that has detected an error and reset its state, also changes its state to the reset state. Hence, an error detection is propagated throughout the graph from node to neighboring node. When a node that is at the reset state is activated, it non-deterministically switches to a state from $C$. Gouda fairness ensures that the graph eventually reaches a correct global configuration.
A node executing Algorithm 1 evaluates the guards in order. If the first guard is enabled, the node performs the corresponding command and deactivates without evaluating the second guard.
Input: The certification scheme $(C,R)$ and a certificate assignment $x_{C}:V\rightarrow C$
1exEvery time node $v$ is activated:
/* states from $C$ are always chosen non-deterministically */
S1 : v.state $\neq\bot$ $\land$ $((B_{1}(v),x_{C}|_{N[v]})\notin R$ $\lor$ $\exists u\in N(v):$ u.state = $\bot)$ $\rightarrow$ v.state $\leftarrow\bot$
S2 : v.state = $\bot$ $\rightarrow$ v.state $\leftarrow$ state from $C$
Algorithm 1 General Certificate to Self-Stabilizing Algorithm
Theorem 5.1.
Let $(C,R)$ be a certification scheme satisfying the Existence and Soundness conditins for some task $\mathcal{T}$ in a connected graph $G$. Starting from any arbitrary configuration, any Gouda fair execution of Algorithm 1 on $G$ stabilizes to a correct final configuration.
Proof 5.2.
Let $G^{*}$ be a configuration where no node satisfies the guard of S1 or of S2, that is, $G^{*}$ is a final correct configuration. By definition, $\mathcal{T}$ is solved in $G^{*}$, since $(C,R)$ is a certification scheme for $\mathcal{T}$. Suppose for contradiction that there exists a configuration $G^{\prime}\neq G^{*}$ that is infinitely often reachable during the execution of Algorithm 1. Since $G^{*}$ is final, it cannot be reachable from $G^{\prime}$. As $G^{\prime}$ is not final, either $G^{\prime}=G_{\bot}$ (i.e., all nodes are in state $\bot$) or the guard of S1 is enabled for some node. In the former case, $G_{\bot}$ trivially appears infinitely often. In the latter, by S1, there exists an activation sequence that leads to $G_{\bot}$. By Gouda fairness, for every a configuration that appears infinitely often, every reachable configuration also appears infinitely often. Hence, in either case $G_{\bot}$ appears infinitely often. Due to S2, once all nodes in $G_{\bot}$ are activated, the configuration becomes $G^{*}$, contradiction. Thus, no configuration $G^{\prime}\neq G^{*}$ can appear infinitely often.
A self–stabilizing algorithm can be obtained by combining the respective certification scheme of each of the previous sections with Algorithm 1.
6 Discussion and Open Problems
In Sections˜3 and 4 we gave certification schemes for leader election, with constant size per edge. A natural direction for further research is studying whether our certification scheme from Section˜4 also works for dismantlable graphs. In particular, only the proof of ˜2.2 depends on the graph being $K_{4}$-free. However, removing this restriction has been very challenging. If this is not possible, another direction is considering bridged or Helly graphs, which are dismantlable but enjoy interesting properties that could be useful for addressing our problem. Alternatively, if our certification scheme can be proven to also work on dismantlable graphs, it is worth exploring other relevant classes. For example, [fieux2020hierarchy] gives a hierarchy on classes (based also on the work of [mazurkiewicz2007locally]) that have dismantling orderings, called s-dismantlable, where 1-dismantlable graphs correspond to the dismantlable graphs defined in Section 4. Informally, for $s>1$, [fieux2020hierarchy] defines a node to be $s+1$ dismantlable if the graph induced by its open neighborhood is $s$-dismantlable and a graph to be $s$-dismantlable if it is reducible to a single vertex by successive deletions of $s$-dismantlable vertices. These directions for future research are summarized in Figure˜4.
Figure 4: A graph class hierarchy, where each arrow represents an inclusion relation. The classes highlighted in green are an easy extension of [chalopin2024stabilising] and those in yellow are considered in this paper for leader election certification with constant memory per edge.
Another potential direction for future research is exploring whether it is possible to automatically transform a certificate into a self-stabilizing algorithm with a weaker scheduler, as defined in the taxonomy of [dubois2011taxonomy], even using non-constant memory. For example, the transformation from [BlinFP14], which uses logarithmic memory but works even with an unfair scheduler, contrasts with our algorithm, which uses constant memory but requires a Gouda fair scheduler. These two results suggest that there may be a trade-off between the strength of the scheduler and the memory requirements that has not yet been fully explored.